OpenStudy (anonymous):

plz help me!!!!!! Simplify 5f + (-3)f + 14. A) 8f + 14 B) f + 14 C) 14 - 2f D) 2f + 14

OpenStudy (nincompoop):

\[5f+(-3)f+14=5f-3f+14\] can you simplify further from here?

OpenStudy (anonymous):

i will atemt to!

OpenStudy (nincompoop):

\[5f-3f=?\]

OpenStudy (anonymous):

how do u simplify? they never really didi teach me to were i understand?

OpenStudy (anonymous):

\[apples ^{2}\]

OpenStudy (anonymous):

so it would be D?

OpenStudy (nincompoop):

yes it is D

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

\[\mathbb{N} \theta \rho \mathbb{R} \theta \beta \iota \epsilon \Omega \Omega \]